how do you solve the problem 3(x-2y)=-12(x+2y)?
Answer by mmm4444bot(95)

Hello There:
We cannot find numeric values for x and y because we only have one equation.
When there are two variables, we need two equations.
If, by chance, you are supposed to solve the equation for y, then you would end up with the slope-intercept form of this linear equation.
You didn't ask any questions or tell us what you're supposed to do with this equation, so I'll just guess that you're supposed to solve for y.
3*(x - 2*y) = -12*(x + 2*y)
Start by using the distributive property to get rid of the parentheses. In other words, both the x and the -2*y get multiplied by 3, and both the x and the 2*y get multiplied by -12.
3*x - 6*y = -12*x - 24*y
Combine like terms. Add 24*y to both sides, and subtract 3*x from both sides.
18*y = -15*x
Solve for y by dividing both sides by 18.
y = (-15/18)*x
y = (-5/6)*x
The graph of this equation is a straight line passing through the origin. Since the slope is a drop of 5 units for every 6 units moved to the right, the point (6, -5) is also on the graph. Plot these two points, and draw a line through them to get the graph.
I hope this is the information that you were looking for. Next time, please consider posting some questions or instructions along with your problem.
~ Mark